INFRARED SPECTRA OF K^{+}(TRYPTAMINE)(H_{2}O)_{n=1−4} AND K^{+}(TRYPTAMINE)(H_{2}O)_{n=0−2}Ar

\underline{AMY L. NICELY} AND JAMES M. LISY, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

A balance of competing electrostatic and hydrogen bonding interactions directs the structure of hydrated gas-phase cluster ions. In K^{+}(Tryptamine) cluster ions, a favorable electrostatic interaction between the potassium cation and the tryptamine NH_{2} lone pair stabilizes the high-energy Gph(in) and Gpy(in) conformers of neutral tryptamine. Previous studies of Tryptamine(H_{2}O)_{n} clusters indicate that the hydrating water molecules stabilize the neutral minimum energy Gpy(out) tryptamine conformer. In this scheme, the first water molecule interacts directly with the NH_{2} lone pair and is located to the side of the tryptamine monomer. By incorporating a potassium cation, however, the minimum energy tryptamine⋯water configuration is disrupted in order to maximize the electrostatic interactions with the cation, shifting so that the tryptamine⋯water interaction includes a π-hydrogen bond between the water and the phenyl ring of tryptamine. The infrared photodissociation spectra of K^{+}(Tryptamine)(H_{2}O)_{n=1−4} and K^{+}(Tryptamine)(H_{2}O)_{n=0−2}Ar will be presented along with parallel ab initio and thermodynamics calculations to assist with the identification of the isomers present in each experiment.